OpenStudy (anonymous):

A boat traveled 280 miles downstream and back. The trip downstream took 7 hours. Trip back took 14 hours. What is the speed of the boat in still water? What is the speed of the current? Please help answer this question. I need to learn how to do this. So teach me every step and thank you. ^.^

OpenStudy (anonymous):

plz hhelp

OpenStudy (kropot72):

Assuming 280 miles is the total distance travelled: Let b = boat speed in still water Let c = current speed. For the downstream trip the speed is b + c. In 7 hours at the speed of (b + c) mph the boat travels 140 miles. 7(b + c) = 140 .............(1) For the upstream trip the speed is b - c. In 14 hours at the speed of (b - c) mph the boat travels 140 miles. 14(b - c) = 140 ............(2) The left hand sides of equations (1) and (2) are equal. Therefore we can write 7b + 7c = 14b - 14c ...........(3) Rearranging equation (3) we get 21c = 7b c = b/3 .......................(4) The value for c obtained in equation (4) should now be substituted into equation (1) which can then be solved to find the value of b. @mariannu Do you follow all of this?

OpenStudy (kropot72):

@Luis_Rivera So you think a total of 560 miles was travelled, 280 miles each way.

OpenStudy (anonymous):

helpful thank you,

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

The word " total" should not did you lead you to believe that you have to split the distances or add them. The distance going is the same as the distance coming, which is 280. So both equation = 280 Irs as simple as that. The asker can verify a solution at any moment.